OpenStudy (anonymous):

A complex one. Integrate (3 + 2cosx) / (2 + 3cosx)^2

OpenStudy (shubhamsrg):

write 3 as 3cos^2x + 3sin^2x and see if you can make it out further

OpenStudy (loser66):

then, partial fraction to solve the integral,

OpenStudy (shubhamsrg):

anyone getting my method? :P

OpenStudy (anonymous):

I don't think there is any ray of hope in it :P

OpenStudy (anonymous):

i need some time.

OpenStudy (shubhamsrg):

hmm, after expressing it as 3cos^2x + 3sin^2x , take cosx common. what i mean to say is : numerator = cosx(3cosx + 2) + 3sin^2x now you see something building up here? :|

OpenStudy (anonymous):

yeah yeah looks like its progressing

OpenStudy (shubhamsrg):

go on! :P try further

OpenStudy (anonymous):

no its not working

OpenStudy (shubhamsrg):

what if I represent numerator like this : (2+ 3cosx)* d(sinx) - sinx * d(2 + 3cosx)

OpenStudy (anonymous):

how ??

OpenStudy (shubhamsrg):

cosx(3cosx + 2) + 3sin^2x = (2+ 3cosx)* d(sinx) - sinx * d(2 + 3cosx) here I am using the fact that derivative of sinx is cosx and of cosx is -sinx

OpenStudy (anonymous):

ok so wats the benefit for using this ?

OpenStudy (shubhamsrg):

try to compare the whole expression with the quotient rule (of derivation)

OpenStudy (anonymous):

ok so it becomes ( sinx )/(2+3cosx) + C ??

OpenStudy (shubhamsrg):

thats right ^_^

OpenStudy (anonymous):

WOW TRULY GENIUS !! HATS OFF BRO !

OpenStudy (shubhamsrg):

a bit of exaggeration but fine, I'll take it :P thank you

OpenStudy (anonymous):

good one!

OpenStudy (anonymous):

One more ?? It's the exclusive integral calculus for Jee main & advanced

OpenStudy (shubhamsrg):

not too sure if I'll be able to pull off this one :3